Fermi National Laboratory
 

Accelerator Update for October 16, 2007 - October 17, 2007

 

Tuesday October 16

The day shift began with the accelerator complex in startup.

Crew Chief Summary: Experts continued with the Tevatron (TeV) power supply dump testing. The TeV quenched twice (once at sector E1 and once at E4). The dump testing concluded after the E4 quench. The repaired D2 power supply was put back in the circuit, but it did not work properly and was taken back out. Beam was injected and circulated in the TeV (without quenching!) – Experts conducted beam studies. SciBooNE / MiniBooNE took beam when it was available. The Antiproton Source (Pbar) conducted reverse proton studies and then established stacking.

Wednesday October 17

The midnight shift began with the accelerator complex in startup.

Crew Chief Summary: TeV experts continued to conduct studies until a Linac RF station (LRF2) went down at 6:30 AM. Pbar experts conducted stacking studies. Recycler conducted proton studies. SciBooNE / MiniBooNE took beam when it was available.

The Plans for Wednesday

The plans for today are to fix LRF2. Problems with the LRF2 modulator interlocks are holding the station off; there can be no beam until this is fixed. Booster, Main Injector (MI), SciBooNE / MiniBooNE, TeV, and Recycler are all ready for beam. The NuMI and SwitchYard (SY) 120 areas are still in shutdown mode. All machines will continue their commissioning efforts.

Machine Summaries

Linac: LRF2 went down and technicians were working on it. They replaced the Linac timing triggers to help resolve their energy problems. They discovered an intertank phase problem that they need 30 minutes or so to tune; this might be contributing to the energy variation problems.

Booster: They would like to tune their new correctors magnet, but can’t due to the Linac energy problems. Experts said that they could deliver about 3E12 of beam.

Main Injector: Yesterday they accessed their tunnel and looked for air leaks, but couldn’t find any major problems. Mechanical techs put another air compressor to help overcome the leaking; this will need further investigation. They confirmed that they would have to remove shielding to repair an LCW leak. Their search for a ground fault stopped when the ground fault went away. They sent beam to Pbar for stacking.

Antiproton Source: Mechanical techs flushed the IB7 magnet and this seemed to clear its overheating problem. They also worked on their shunts while they were in on access. After the access they conducted Pbar Debuncher orbit studies and worked on cooling. Some RF problems came up that they said they would be fixed today. They would like to work on stack phasing today.

Recycler: Experts conducted proton studies. They are ready for antiprotons and need them for phasing.

Tevatron: EE support experts completed their testing. TeV experts successfully injected beam into the TeV and the orbits looked good. They need to access the tunnels to fix some monitoring equipment. They would like to accelerate beam tonight.

Cryo: The TeV sector F4 problem, which the cryo system techs fixed, was due to a leak and contamination in the vacuum system. They would need 3-4 hours to fix the D4 wet engine.

Controls: Experts are still trying to deal with the alarms problem. Whenever a large amount of alarms post at the same time the system goes down.

D0: Experts enabled the SVX beam permit. They still need 8-10 hours to open their detector and replace a bad fuse.

CDF: They are ready for beam. They only have a few small jobs that would require an access.

FESS: They found a problem with TeV sector D4 motor control center, which they will work on today.

Mechanical: The Meson M01 area has a water skid with a failing LCW pump.

MiniBooNE/SciBooNE: An expert reported that they had a problem with their autotune program that required a restart to fix.

SwitchYard 120: They are still in shutdown. A new experiment will start up on November 7 th.

NuMI: They are still in shutdown. An expert reminded everyone that they still needed to test the HV101 power supply, which is a critical device.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/17/2007    email Fermilab